Find the amount (future value) of an ordinary annuity of $1600 per semiannual period for 7 years at 8% per year compounded semiannually. (Round your answer to the nearest cent.)
Answer by mananth(16949)

Principal P = 1600
Amount= A
years=n 7
compounded 2 times a year
Rate = 8.00 0.04
Amount = P*((n+r)/n)^n

Amount =1600*(1 +0.04)^ 7*2
Amount =1600*( 1+0.04)^14
1600*(1.04)^14
Amount = 2770.68